RC Meeting and Run Plan, Sunday, December 15th 2019

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 4:15 PM, 2nd floor Counting House Meeting Room

  • Bluejeans remote connection information:
 To join the Meeting:
 To join via Room System:
 Video Conferencing System: bjn.vc -or-
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links

Current Status

  • Online Analysis Update and instructions???
  • Priority: Target Work
  • Working on maximizing long. laser polarization
  • Techs and Radcon will be here at 7am to close up enclosure (~6hours)
  • Spectrometer Magnets Off (Alarms Masked)[1]
  • PNMR FID signal can't be found. System can be troubleshooted in CH. [2]
  • 3 NMRs are working
  • EPR is working [3] but noisy [4]
  • Ramp rate lock instructions are here and added to checklist[5]
  • Temporary Laser instructions can be found here [6]
  • Techs pumped down alarming transfer tube yesterday [7]

Current Run Plan

  • Run Plan
  • Energy measurement at 5-pass [8]
  • Day by Day Options [9] Can we add Moller shift Monday Swing if we have 5-pass (For accidental study)?

Notices in Effect for Shift Crew

  • Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
  • Magnet policies
    • HMS Q1 must be ramped down in 50A steps in all cases
    • Do not ramp HMS Q2 to zero, only go to 10A and then switch off power.
  • There IS a lock available for the ramp rate. I will post instructions on how to check if it was on.
  • Inform the RC of any access that the Target Team requires. The RC will coordinate with RadCon to make everything run smoothly.
  • The operators have been somewhat careful with our beam orbit locks. Please make sure that the orbit locks are enabled and the beam does not drift from its nominal position.

Current Issues

  • Long. Laser
  • SHMS Dipole Power Supply
  • HMS Dipole LN2 overfilling (Will check tomorrow

Ongoing Work in the Hall

  • Carbon Foils (Done)
  • Checking Optical Fibers (Done)
  • Maximizing Long. Laser Polarization (In Progress)
  • Resolder EPR connection coils (Done) [10]
  • Bunker Fan (Done) [11]
  • Replace fan in optics table (Done?)
  • Tranverse NMR pick-up coil (Done)
  • PNMR RF switch check (Done)
  • Transverse Laser Alignment (Done) [12]
  • Fiber Check [13]

Planned Accelerator Configuration Changes

  • Change to one pass Monday or Tuesday

Special Requests

  • Keep target expert schedule filled out on whiteboard